The area enclosed between the curve y = loge (x + e) and the coordinate axes is ?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: 1
Area ‘A’:
A = \int_{1-e}^{0}ln(x+e)dx
A = \int_{1-e}^{0}x^{0}ln(x+e)dx
A = (x.ln(x+e))_{1-e}^{0} -\int_{1-e}^{0} \frac{x}{x+e}dx
A = -\int_{1-e}^{0} \frac{x+e-e}{x+e}dx
A = \int_{1-e}^{0} (\frac{e}{x+e}-1)dx
A = (e.ln(x+e) -x)_{1-e}^{0} = 1
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty